### v1

@@ II-C Reliability Analysis
This yields two BPSK symbols. Hard decisions $\hat{\tilde{b}}_{1}=\operatorname{sgn}(r_{\mathrm{I}})$ and $\hat{\tilde{b}}_{0}=\operatorname{sgn}(r_{\mathrm{Q}})$, followed by $\hat{b}_{k}=(1-\hat{\tilde{b}}_{k})/2$, $k\in\{0,1\}$, extract the two bits. Since AWGN is symmetric, this hard-decision scheme induces two binary symmetric channels (BSCs) with probability of error:
@@
The ECC structure, which can aid Willie in detection, is eliminated by applying a $2n_{t}$-bit one-time-pad $\vec{s}$ to the encoder output written in binary. Each bit in $\vec{s}$ is selected by Alice and Bob equiprobably at random, resulting in output distribution $\Pr(x=-1)=\Pr(x=1)=\frac{1}{2}$. Pre-shared secret includes $\vec{t}$ and $\vec{s}$.
@@ II-D Hypothesis Testing and Covertness
Willie observes sequence $\vec{w}\in\mathbb{C}^{n}$ and has to decide if Alice is transmitting. Unlike Bob, he does not have access to $\vec{t}$ and $\vec{s}$, but he knows the exact transmission start time, and pulse boundaries. Additionally, we assume that he knows the channel conditions, and other details of transceiver design (including $\vec{c}$ and $\alpha$). He performs a binary hypothesis test between hypotheses $H_{0}$ (no transmission) and $H_{1}$ (transmission). Let distributions $P_{0}^{n}$ and $P_{1}^{n}$ and associated density functions $p_{0}^{n}(\vec{w})$ and $p_{1}^{n}(\vec{w})$ describe the statistics of Willie’s output $\vec{w}$ when Alice is silent ($H_{0}$) and transmitting ($H_{1}$). Assuming non-informative priors [^1] Accounting for arbitrary priors is discussed in [38]. $\Pr(H_{0})=\Pr(H_{1})=\frac{1}{2}$, Willie’s probability of error for an optimal detection scheme is [43, Th. 13.1.1]:
@@
where $\mathcal{V}_{T}(P_{0}^{n},P_{1}^{n})\triangleq\frac{1}{2}\int_{\mathbb{V}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}|p_{0}^{n}(\vec{w})-p_{1}^{n}(\vec{w})|$ is the total variation distance between $P_{0}^{n}$ and $P_{1}^{n}$. We say that the transmission is $\delta$-covert if $p_{e}^{(w)}\geq\frac{1}{2}-\delta$. Total variation distance is mathematically unwieldy, hence, in covert communication literature [1, 2, 9, 10] we often employ Pinsker’s inequality [8, Lemma 11.6.1] to lower bound
@@
where $D(P_{0}^{n}\|P_{1}^{n})\triangleq\int_{\mathbb{C}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}p_{0}^{n}(\vec{w})\log_{2}\frac{p_{0}^{n}(\vec{w})}{p_{1}^{n}(\vec{w})}$ is the relative entropy of $P_{0}^{n}$ and $P_{1}^{n}$. Thus, instead of (9), it is standard to use (10), and assert that any scheme is $\delta$-covert if $D(P_{0}^{n},P_{1}^{n})\leq\delta_{\rm RE}=8\delta^{2}$. Ultimately, satisfying this relative entropy constraint for a given $\delta$ requires limiting our average output power. While relative entropy between our distributions can be readily bounded, Pinsker’s inequality is loose for many distribution pairs. Thus, we seek a tighter bound than (10), which will allow Alice more average power throughput while still satisfying (9). Inspired by recent work [27] that characterized covert communication over a quantum channel model with fidelity constraint, we employ the classical analog of the fidelity constraint to instead bound (9) via the Cauchy-Schwarz inequality as follows:
@@
When not silent, Alice inputs QPSK symbol $x\in\mathcal{X}$ in each pulse slot according to Section II-B. Then, Willie receives $\vec{w}_{p}(x)=h_{w}\vec{u}(x)+\vec{z}^{(w)}$ for $\vec{u}(x)=\left(e^{j\theta}\vec{c}_{p},e^{j\theta}\vec{c}_{q}(x)\right)$, per the model in Fig. 1. Let $r_{p/q}=\frac{\|\vec{c}_{p}\|}{\|\vec{c}_{q}\|}$ be the ratio of magnitudes of the pilot and QPSK pulses, and $c_{q}=\|\vec{c}_{q}\|$ be the magnitude of the QPSK pulse. Appendix A shows that we may ensure $\delta$-covertness of the transmission scheme by setting
$\alpha_{n}$ $=\frac{4\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{2\log\left(\frac{1}{1-2\delta^{2}}\right)}{n_{p}(1+r_{p/q}^{4})}}$ (12)
$=\left(\mathrm{SNR}\sqrt{1+r_{p/q}^{4}}\right)^{-1}\sqrt{\frac{32\log\left(\frac{1}{1-2\delta^{2}}\right)}{n_{p}}},$ (13)
@@
III Experiment Implementation and Results
@@ III-A System Configuration
We perform our experiments on ORBIT, an open-access radio grid testbed [32]. Our system is depicted in Fig. 4. We connect four Ettus universal software radio peripheral (USRP) X310 SDRs fitted with UBX 160 – corresponding to transmitter Alice, receiver Bob, adversary Willie, and an AWGN generator – into a shared-medium RF star network topology via coaxial cables. This isolates our experiment from others on ORBIT, enables control over the environment, and allows repeatability. We defer a wireless experiment with an organic noise source to future work. The radios are positioned so that Bob and Willie experience approximately equal attenuation from both the AWGN generator and Alice, corresponding to an adverse scenario of Willie being close to Bob. We utilize a single 12.5 MHz channel centered at $f_{c}=915$ MHz.
@@
Each radio has a dedicated enhanced small form-factor pluggable (SFP+) cable connecting it to a high-bandwidth router, and the router to a dedicated control node on the ORBIT grid [33]. SFP+ ports operate in 10 Gbps mode to support the maximum transmission unit (MTU) size of 8 kB. This prevents packet drops from the internal radio buffer overflow during an experiment. We note that even a single packet drop causes catastrophic misalignment of the covert symbols within the transmission, rendering useless an entire experimental trial. ORBIT provides a network [33] connecting the control nodes to network attached storage (NAS) and to other nodes forming a compute cluster for processing the collected data.
@@
Figure 4: Covert communication experiment on ORBIT. Four Ettus USRP X310 radios – Alice (Tx), Bob (Rx), Willie (warden), and a broadband noise source – are linked by coaxial cables in a star topology using Mini-Circuits ZFSC-2-10G splitters/combiners. Tx-to-Rx and Tx-to-Tx path losses are 50 dB and 65 dB, respectively. All radios operate at $f_{c}=915\,\text{MHz}$ with Alice, Bob, and Willie using a DAC/ADC sampling rate of $f_{s}=12.5\times 10^{6}$ samples/s and the noise generator $f_{s}=25\times 10^{6}$ samples/s. Alice, Bob, and Willie apply 0 dB Tx/Rx gain while the noise generator applies 20 dB gain. Each X310 connects over a 10 Gb/s SFP+ link to a dedicated control node (Intel Xeon E5-2640, 20 cores); Alice’s node orchestrates the experiment via TCP messages to the other nodes and an eleven-node compute cluster of the same machines that performs real-time processing while a 2 TB network attached storage (NAS), mounted via NFS v4.2, provides shared buffer.
@@
The radios’ internal clocks are synchronized using an Ettus OctoClock, which provides low-jitter pulse-per-second (PPS) and 10 MHz reference signals, allowing the radios to maintain a constant phase offset between each other [44]. We note that a centralized clock source is merely an experimental convenience. In practice, Alice and Bob can synchronize independently to any stable clock source, such as an atomic clock, prior to transmission. [^2] In a separate experiment in our lab at the University of Arizona, we confirmed this using a Stanford Research Systems FS725 10 MHz Rubidium Frequency Standard connected to Ettus USRP N210 radios.
Figure 5: Data-packet structure (durations in samples). A 15 400-sample preamble is created by concatenating five 13-bit Barker sequences, BPSK-modulating them, zero-padding each to 200 samples, and convolving the result with a 2 401-sample root-raised-cosine filter. The subsequent $T$-second, $T\times f_{s}$-sample, segments encode the two parts of the experimental trial. In the Alice-on segment, each shaded pulse slot begins with a strictly positive, real-valued Gaussian pilot pulse (for phase reference), immediately followed by a Gaussian-shaped, QPSK-modulated data pulse carrying randomly generated information bits. The indices of the occupied slots are drawn uniformly at random and stored in $\vec{t}$ (see Section II-C). In the subsequent Alice-off segment, she remains silent.
@@
Alice’s node generates fresh pre-shared secret and message vectors $\vec{t}$ and $\vec{x}_{2n_{t}}$ for each experimental trial. Bob and Willies’ radios sample the channel continuously while Alice supplements the RF trial packet described in Section III-D with TCP control messages over the SFP+ connections that mark its precise start and stop times. At the end of each trial, control nodes write their data to the network-attached storage (NAS): Alice’s node stores $\vec{t}$ and $\vec{x}_{2n_{t}}$, whereas Bob and Willies’ nodes write their raw IQ samples, padded with brief pre- and post-buffers. Alice’s node then alerts the compute cluster, which analyzes the new trial, and deletes its data on completion. Acting as a rolling buffer, the NAS keeps heavy data traffic off the radio-control links and holds disk usage well below its 2 TB capacity. The Ettus USRP library and USRP hardware driver (UHD), [45, 46] are used to interact with the radios.
III-B Spectral Footprint of Alice’s Transmissions
Our analysis in Section II-D assumes that Alice’s signal energy is confined to a single $12.5$ MHz channel. In practice, this is often violated: time-limited signals are unbounded in frequency domain, and transmitter hardware non-idealities can generate strong spectral artifacts.
@@ III-B Spectral Footprint of Alice’s Transmissions
To characterize the emitted spectrum, we capture Alice’s waveform with a Tektronix RSA5115B real-time spectrum analyzer positioned at Willie’s location so that it records exactly what the warden observes. The analyzer is centered at $915$ MHz, samples at $2\times 10^{8}$ samples/s, and the external AWGN generator remains disabled throughout these spectral measurements. Alice transmits the shortest Gaussian pulses permitted by the system, $n_{s}^{(p)}=n_{s}^{(q)}=6$, at an amplitude of $m_{p}=0.12351232,m_{q}=1.3313578$.
@@
Figure 6: Measured spectrum at Willie’s receiver. The Gaussian main lobe, expected to be centered at 915 MHz, is displaced by approximately +300+300 Hz, indicating a small carrier–frequency offset.
@@
Figure 7a overlays the pilot–data pulse pair at the transmitter with the corresponding waveform captured by the spectrum analyzer; the traces align closely, confirming that the temporal shape is maintained. Figure 7b presents the received spectrum over a wide frequency span and includes an inset which zooms to the carrier band observable at the transmitter. The transmit and receive spectra are overlaid in the inset and coincide nearly identically, confirming that the Gaussian envelope and its 99.9% energy bandwidth are maintained. The full-span view further shows that essentially all energy is confined to this zoomed region; only a few narrow spikes appear in the immediate vicinity of the band edges, and with a single capture it is unclear whether these are random artifacts or a repeatable spectral feature
Farther out in frequency, Figure 8a shows two pronounced spurs at the third and fifth harmonics of the carrier. Odd harmonics typically arise when a nominally sinusoidal signal develops a more square-like waveform, which in a transmitter can happen either through power-amplifier compression or through non-sinusoidal drive in the local oscillator (LO) chain. Neither explanation fits neatly here: the measurement is made at 0 dB hardware gain, well below the power-amplifier’s compression point, and the UBX-160 LO path is followed by a 2.2 GHz low-pass section that should strongly attenuate components at 2.75 GHz and 4.58 GHz. The precise mechanism that leaves these third- and fifth-harmonic spurs above the noise floor is therefore still under investigation.
@@
Figure 7: Carrier-band verification of the 12-sample pulse pair transmitted by Alice at 0 dB gain. Each pair comprises a 6-sample Gaussian pilot followed immediately by a 6-sample Gaussian QPSK-data pulse. The left panel overlays the time-domain waveform at the transmitter (TX) and at a spectrum-analyzer (RX) placed at Willie, while the right panel shows the RX spectra, with insets overlaying the TX and RX spectra in the carrier band.
(a) Repeated pilot–data pulse pair
@@
Figure 8: Spectra at the first five harmonics, measured at Willie with Alice’s transmitter gain set to 0 dB. Left: when Alice continuously repeats the pilot–data pulse pair. Right: when the baseband input is forced to all zeros (silent mode). The ordinate scale is re-ranged from one harmonic panel to the next to keep weak spurs visible.
To verify that Alice’s silent mode is truly quiet, we transmit an all-zero waveform. Even with an all-zero signal, the spectrum shows a small spur at the carrier and faint harmonics (Fig. 8b), behavior attributable to LO leakage.
The two spectral artifacts most relevant to covertness at the settings used in our experiments are (i) the power that spills into odd harmonics and (ii) the residual carrier spur that appears when Alice is commanded to remain silent. The former could be mitigated by inserting an external low-pass filter at Alice’s RF output, while the latter may be reduced through a careful power calibration of the X310 transmit chain. While making these improvements is challenging on a shared-testbed hardware with limited on-site access, we will pursue this in the follow-on work.

### v2

@@ II-C Reliability Analysis
This yields two BPSK symbols. Hard decisions $\hat{\tilde{b}}_{1}=\operatorname{sgn}(r_{\mathrm{I}})$ and $\hat{\tilde{b}}_{0}=\operatorname{sgn}(r_{\mathrm{Q}})$, followed by $\hat{b}_{k}=(1-\hat{\tilde{b}}_{k})/2$, $k\in\{0,1\}$, extract the two bits. Since AWGN is symmetric, this hard-decision scheme induces two binary symmetric channels (BSCs) with a probability of error:
@@
The ECC structure, which can aid Willie in detection, is eliminated by applying a $2n_{t}$-bit one-time pad $\vec{s}$ to the binary encoder output. Each bit in $\vec{s}$ is selected by Alice and Bob equiprobably at random, resulting in output distribution $\Pr(x=-1)=\Pr(x=1)=\frac{1}{2}$. Pre-shared secret includes $\vec{t}$ and $\vec{s}$.
@@ II-D Hypothesis Testing and Covertness
Willie observes the sequence $\vec{w}\in\mathbb{C}^{n}$ and must decide whether Alice is transmitting. Unlike Bob, he does not have access to $\vec{t}$ and $\vec{s}$, but he knows the exact transmission start time and pulse boundaries. That is, Willie can access Alice and Bob’s shared time, which we model in the experiment by using a common clock for all parties. Additionally, we assume that he knows the state of his channel from Alice, and other details of Alice and Bob’s transceiver design (including $\vec{c}_{p}$, $\vec{c}_{q}$, and $\alpha_{n}$). The stringency of these assumptions ensures that the system is secure against the most capable adversary. We defer studying the impact of relaxing these to future work.
Willie performs a test between binary hypotheses $H_{0}$ (silence) and $H_{1}$ (transmission). Let $P_{\text{FA}}=P(\text{choose~}H_{1}|H_{0})$ and $P_{\text{MD}}=P(\text{choose~}H_{0}|H_{1})$ denote the probabilities of false alarm and missed detection, respectively. Assuming non-informative priors [^1] Accounting for arbitrary priors is discussed in [40]. $\Pr(H_{0})=\Pr(H_{1})=\frac{1}{2}$, Willie’s probability of error is:
$p_{e}^{(w)}$ $\triangleq\frac{1}{2}\left(P_{\text{FA}}+P_{\text{MD}}\right).$ (9)
Let distributions $P_{0}^{n}$ and $P_{1}^{n}$ and associated density functions $p_{0}^{n}(\vec{w})$ and $p_{1}^{n}(\vec{w})$ describe the statistics of Willie’s output $\vec{w}$ when Alice is silent ($H_{0}$) and transmitting ($H_{1}$). An optimal detection scheme yields [29, Th. 15.1.1]:
@@
where $\mathcal{V}_{T}(P_{0}^{n},P_{1}^{n})\triangleq\frac{1}{2}\int_{\mathbb{V}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}|p_{0}^{n}(\vec{w})-p_{1}^{n}(\vec{w})|$ is the total variation distance between $P_{0}^{n}$ and $P_{1}^{n}$. We say that the transmission is $\delta$-covert if $p_{e}^{(w)}\geq\frac{1}{2}-\delta$. Total variation distance is mathematically unwieldy, hence, in covert communication literature [9, 10, 14, 47] we often employ Pinsker’s inequality [19, Lemma 11.6.1] to lower bound
@@
where $D(P_{0}^{n}\|P_{1}^{n})\triangleq\int_{\mathbb{C}^{n}}\mathop{}\!\mathrm{d}^{2}\vec{w}p_{0}^{n}(\vec{w})\log_{2}\frac{p_{0}^{n}(\vec{w})}{p_{1}^{n}(\vec{w})}$ is the relative entropy of $P_{0}^{n}$ and $P_{1}^{n}$. Thus, instead of (10), one can use (11), and assert that any scheme is $\delta$-covert if $D(P_{0}^{n}\|P_{1}^{n})\leq\delta_{\rm RE}=8\delta^{2}$. Ultimately, satisfying this relative entropy constraint for a given $\delta$ requires limiting Alice’s average output power. While relative entropy between our distributions can be readily bounded, Pinsker’s inequality is loose for many distribution pairs. Thus, we seek a tighter bound than (11) to increase Alice’s output power while satisfying (10). Inspired by recent work [49] that characterized covert communication over a quantum channel using quantum fidelity, we employ the classical analog of quantum fidelity to instead bound (10) [29, Th. 15.1.2] via the Cauchy-Schwarz inequality:
@@
Recall from Section II-B that, when transmitting, Alice inputs QPSK symbol $x\in\mathcal{X}$ in each pulse slot. Then, Willie receives $\vec{w}_{p}(x)=h_{w}\vec{u}(x)+\vec{z}^{(w)}$ for $\vec{u}(x)=\left(e^{j\theta}\vec{c}_{p},e^{j\theta}\vec{c}_{q}(x)\right)$, per the model in Fig. 1. Let $r_{p/q}=\frac{\|\vec{c}_{p}\|}{\|\vec{c}_{q}\|}$ be the ratio of magnitudes of the pilot and QPSK pulses, and $c_{q}=\|\vec{c}_{q}\|$ be the magnitude of the QPSK pulse. Appendix A shows that we ensure $\delta$-covertness of the transmission scheme by setting
$\alpha_{n}$ $\leq\frac{4\sqrt{2}\sigma_{w}^{2}}{a_{w}^{2}c_{q}^{2}}\sqrt{\frac{-\log\left(1-2\delta^{2}\right)}{n_{p}(1+r_{p/q}^{4})}}$
$=\left(\mathrm{SNR}\sqrt{1+r_{p/q}^{4}}\right)^{-1}\frac{4\sqrt{2}}{\sqrt{n_{p}}}\sqrt{-\log\left(1-2\delta^{2}\right)},$ (13)
@@
II-E Willie’s Detectors
The preceding analysis abstracts away detector structures available to Willie. We discuss them here, deferring their analysis to Appendices C and D.
1 Optimal Detector
For known densities $p_{0}^{n}(\vec{w})$ and $p_{1}^{n}(\vec{w})$, a likelihood ratio test (LRT) is optimal and achieves $p_{e}^{(w)}$ in (10) with equality as $n\to\infty$ [29, Th. 15.1.1]. LRT also minimizes $P_{\text{MD}}$ for a constant $P_{\text{FA}}$, per Neyman-Pearson lemma [29, Th. 3.2.1]. Here, $p_{0}^{n}(\vec{w})=\prod_{i=1}^{n_{p}}p^{(0)}\left(\vec{w}_{i}\right)$ and $p_{1}^{n}(\vec{w})=\prod_{i=1}^{n_{p}}p^{(1)}\left(\vec{w}_{i}\right)$, with $p^{(0)}\left(\vec{w}_{i}\right)$ and $p^{(1)}\left(\vec{w}_{i}\right)$ in (18) and (19), where $\vec{w}_{i}$ are the observations of the $i^{\text{th}}$ pulse slot. Willie accuses Alice if the LRT statistic $L(\vec{w})<0$, where
$L(\vec{w})\triangleq\log\frac{p_{0}^{n}(\vec{w})}{p_{1}^{n}(\vec{w})}=\sum_{i=1}^{n_{p}}L_{i}$ (14)
and $L_{i}=\log\frac{p^{(0)}(\vec{w}_{i})}{p^{(1)}(\vec{w}_{i})}$. Willie computes $L(\vec{w})$ from observations in $\vec{w}$ given his knowledge of Alice and Bob’s system. Unfortunately, Bessel functions in $p^{(1)}\left(\vec{w}_{i}\right)$ complicate the analysis of the resulting detector. However, we expect the QPSK pulse magnitude $c_{q}$ to be small. Employing the Taylor series expansion of $L(\vec{w})$ around $c_{q}=0$ in Appendix C, we simplify the LRT: Willie accuses Alice if the test statistic $S^{\text{(opt)}}(\vec{w})\geq\tau^{\text{(opt)}}$, where $S^{\text{(opt)}}(\vec{w})=\sum_{i=1}^{n_{p}}S_{i}^{\text{(opt)}}$,
$S_{i}^{\text{(opt)}}$ $=\left|\left\langle\vec{c}_{p},\vec{p}_{i}\right\rangle\right|^{2}+\left|\left\langle\vec{c}_{q},\vec{q}_{i}\right\rangle\right|^{2},$ (15)
$\vec{w}_{i}=(\vec{p}_{i},\vec{q}_{i})$ is the decomposition into pilot and QPSK segments, and $\tau^{\text{(opt)}}$ is the threshold discussed later.
2 Power Detector
Per (15), Willie needs synchronization with Alice to use the optimal detector. Radiometer offers a practical alternative: Willie accuses Alice if the total collected power $\left\|\vec{w}\right\|^{2}$ exceeds the threshold $\tau^{\text{(rad)}}$. For convenience, we define the radiometer test statistic $S^{\text{(rad)}}(\vec{w})=\sum_{i=1}^{n_{p}}S_{i}^{\text{(rad)}}$, where $S_{i}^{\text{(rad)}}=\left\|\vec{w}_{i}\right\|^{2}$ is in correspondence to (15).
In Appendix D we employ the Berry-Esseen theorem to analyze the probability of error achievable by these schemes, and derive the thresholds $\tau^{\text{(opt)}}$ and $\tau^{\text{(rad)}}$. Furthermore, since the optimal detector achieves $p_{e}^{(w)}$ in (10), we show in Appendix D-B how to improve $\alpha_{n}$ from (13) to
$\alpha_{n}$ $\leq\left(\mathrm{SNR}\sqrt{1+r_{p/q}^{4}}\right)^{-1}\frac{4\sqrt{2}}{\sqrt{n_{p}}}\sqrt{\pi}\delta.$ (16)
However, we use more conservative (13) in our experiments. [^2] An anonymous referee’s final review comment led to (16). Regrettably, the publication deadline prevented us from using it in our experiments.
III Experiment Implementation
@@ III-A System Configuration
We perform our experiments on COSMOS (formerly ORBIT), an open-access radio grid testbed [33, orbit_grid]. Our system is depicted in Fig. 4. We connect four Ettus universal software radio peripheral (USRP) X310 SDRs, each fitted with a UBX 160 – corresponding to transmitter Alice, receiver Bob, adversary Willie, and an AWGN generator – into a shared-medium RF star network via coaxial cables. This isolates our experiment from others on COSMOS, enables control over the environment, and allows repeatability. We defer a wireless experiment with an organic noise source to future work. Our radios are not power-calibrated. Therefore, we report measurements of signal amplitude and power in arbitrary units, respectively denoted as a.a.u. and a.p.u. The radios are positioned so that Bob and Willie experience approximately equal attenuation from both the AWGN generator and Alice, corresponding to an adverse scenario in which Willie is close to Bob. We utilize a single 12.5 MHz channel centered at $f_{c}=915$ MHz. Both the transmitting (Alice and noise generator) and receiving (Bob and Willie) radios are offset-tuned by 25 MHz.
@@
Each radio has a dedicated enhanced small form-factor pluggable (SFP+) cable connecting it to a high-bandwidth router, and the router to a dedicated control node on the COSMOS grid [orbit_grid]. SFP+ ports operate at 10 Gbps to support a maximum transmission unit (MTU) of 8 kB. This prevents packet drops due to an internal radio buffer overflow during an experiment. We note that even a single packet drop can cause catastrophic misalignment of the covert symbols during transmission, rendering an entire experimental trial useless. COSMOS provides a network [orbit_grid] connecting the control nodes to network attached storage (NAS) and to other nodes forming a compute cluster for processing the collected data.
@@
Figure 4: Covert communication experiment on COSMOS. Four Ettus USRP X310 radios – Alice (Tx), Bob (Rx), Willie (warden), and a broadband noise source – are linked by coaxial cables in a star topology using Mini-Circuits ZFSC-2-10G splitters/combiners. Tx-to-Rx and Tx-to-Tx path losses are 50 dB and 65 dB, respectively. All radios operate at $f_{c}=915\,\text{MHz}$ with Alice, Bob, and Willie using a DAC/ADC sampling rate of $f_{s}=12.5\times 10^{6}$ samples/s and the noise generator $f_{s}=22.5\times 10^{6}$ samples/s. Alice, Bob, and Willie apply 0 dB Tx/Rx gain while the noise generator applies 15 dB gain. Each X310 connects over a 10 Gb/s SFP+ link to a dedicated control node (Intel Xeon E5-2640, 20 cores); Alice’s node orchestrates the experiment via TCP messages to the other nodes and an eleven-node compute cluster of the same machines that performs real-time processing, while a 2 TB network-attached storage (NAS), mounted via NFS v4.2, provides a shared buffer.
@@
The radios’ internal clocks are synchronized using an Ettus OctoClock, which provides low-jitter pulse-per-second (PPS) and 10 MHz reference signals, allowing the radios to maintain a constant phase offset between each other [octoclock_datasheet]. We note that a centralized clock source is merely an experimental convenience. In practice, Alice, Bob, and Willie can synchronize independently to any stable clock source, such as an atomic clock, prior to transmission. [^3] In a separate experiment in our lab at the University of Arizona, we confirmed this using a Stanford Research Systems FS725 10 MHz Rubidium Frequency Standard connected to Ettus USRP N210 radios.
@@
Alice’s node generates fresh pre-shared secret and message vectors $\vec{t}$ and $\vec{x}_{2n_{t}}$ for each experimental trial. Bob and Willie’s radios continuously sample the channel, while Alice supplements the RF trial packet described in Section III-D with TCP control messages over the SFP+ connections that mark its precise start and stop times. At the end of each trial, control nodes write their data to the network-attached storage (NAS): Alice’s node stores $\vec{t}$ and $\vec{x}_{2n_{t}}$, whereas Bob and Willie’s nodes write their raw IQ samples, padded with brief pre- and post-buffers. Alice’s node then alerts the compute cluster, which analyzes the new trial and deletes its data on completion. Acting as a rolling buffer, the NAS keeps heavy data traffic off the radio-control links and holds disk usage well below its 2 TB capacity. The Ettus USRP library and USRP hardware driver (UHD), [uhd_github, uhd_manual] are used to interact with the radios.
III-B Alice’s Spectral Footprint and SDR Non-idealities
Section II-D addresses the fundamental limitations of implementing covert communication on digital transceivers. Here, we identify the constraints specific to our SDR platform.
@@ III-B Alice’s Spectral Footprint and SDR Non-idealities
Our analysis assumes that Alice’s signal energy is confined to a single $12.5$ MHz channel. In practice, this is often violated: time-limited signals are unbounded in the frequency domain, and transmitter hardware non-idealities can generate strong spectral artifacts. To characterize the emitted spectrum, we capture Alice’s waveform with a Tektronix RSA5115B real-time spectrum analyzer positioned at Willie’s location to record his observations exactly. The analyzer is centered at $915$ MHz, samples at $2\times 10^{8}$ samples/s, and the external AWGN generator remains disabled throughout these spectral measurements. Alice transmits the shortest Gaussian pulses permitted by the system, $n_{s}^{(p)}=n_{s}^{(q)}=6$, at an amplitude of $m_{p}=0.12351232,m_{q}=1.3313578$.
@@
Figure 5: Measured spectrum at Willie’s receiver. The Gaussian main lobe, expected to be centered at 915 MHz, is displaced by approximately +300+300 Hz, indicating a small carrier-frequency offset.
@@
Fig. 6LABEL:sub@fig:gaussian_single_pulse_iq overlays the pilot-data pulse pair at the transmitter with the corresponding waveform captured by the spectrum analyzer; the traces align closely, confirming that the temporal shape is maintained. Fig. 6LABEL:sub@fig:gaussian_single_pulse_spectrum presents the received spectrum over a wide frequency span and includes an inset that zooms to the carrier band observable at the transmitter. The transmit and receive spectra are overlaid in the inset and coincide nearly identically, confirming that the Gaussian envelope and its 99.9% energy bandwidth are maintained. The full-span view further shows that essentially all energy is confined to this zoomed region; only a few narrow spikes appear in the immediate vicinity of the band edges, and with a single capture, it is unclear whether these are random artifacts or a repeatable spectral feature.
Farther out in frequency, Fig. 7LABEL:sub@fig:harmonics_tx_0dB shows two pronounced spurs at the third and fifth harmonics of the carrier. Odd harmonics typically arise when a nominally sinusoidal signal develops a more square-like waveform, which in a transmitter can happen either through power-amplifier compression or through non-sinusoidal drive in the local oscillator (LO) chain. Neither explanation fits neatly here: the measurement is made at 0 dB hardware gain, well below the power amplifier’s compression point, and the UBX-160 LO path is followed by a 2.2 GHz low-pass section that should strongly attenuate components at 2.75 GHz and 4.58 GHz. The precise mechanism that leaves these third- and fifth-harmonic spurs above the noise floor, therefore, should be investigated.
@@
Figure 6: Carrier-band verification of the 12-sample pulse pair transmitted by Alice at 0 dB gain. Each pair comprises a 6-sample Gaussian pilot followed immediately by a 6-sample Gaussian QPSK-data pulse. The left panel overlays the time-domain waveforms at the transmitter and at a spectrum analyzer placed at Willie, while the right panel shows the spectrum analyzer output, with insets overlaying the transmitted and received spectra in the carrier band. The amplitude and power transmitted by our uncalibrated radio are in arbitrary amplitude and power units (a.a.u. and a.p.u.).
(a) Repeated pilot-data pulse pair
@@
Figure 7: Spectra at the first five harmonics, measured at Willie with Alice’s transmitter gain set to 0 dB. Left: when Alice continuously repeats the pilot-data pulse pair. Right: when the baseband input is forced to all zeros (silent mode). The ordinate scale is in dB relative to peak transmitted power.
To verify that Alice’s silent mode is truly quiet, we transmit an all-zero waveform. Even with an all-zero signal, the spectrum in Fig. 7LABEL:sub@fig:harmonics_silent_0dB shows a small spur at the carrier and faint harmonics, behavior attributable to LO leakage.
The two spectral artifacts most relevant to covertness at the settings used in our experiments are (i) the power that spills into odd harmonics and (ii) the residual carrier spur that appears when Alice is commanded to remain silent. The former could be mitigated by inserting an external low-pass filter at Alice’s RF output, while the latter may be reduced through a careful power calibration of the X310 transmit chain. Making these improvements is challenging on shared-testbed hardware with limited on-site access. We will pursue this in the follow-on work.
The radio circuitry is always on in our experiments. However, prior theoretical work and the analysis in Section II effectively assume that Alice’s transmitter is completely powered off unless it is transmitting a pulse. In practice, power cycling a radio consumes significant time and may emit unwanted signals. That being said, we will investigate powering it off in the future, as the long intervals between sparse-coded transmissions may permit it. Furthermore, in some practical scenarios, $H_{0}$ can correspond to Alice’s transmitter being always on but transmitting zero, as in our implementation.
Finally, LO power leaks into the baseband due to imperfect isolation between the LO and RF ports. When combined with small frequency offsets of the LO from the desired center frequency $f_{c}$, this leakage manifests as high-amplitude low-frequency noise. To mitigate this, we offset-tune our radios by 25 MHz, shifting the signal of interest away from this DC-centered noise.